OpenStudy (anonymous):

SIMPLIFY THE RATIO?? 4.4 TO 3.6...

OpenStudy (anonymous):

HAS TO BE AS A FRACTION!!!

OpenStudy (anonymous):

Ratio is just same as Divide..

OpenStudy (anonymous):

\[\frac{4.4}{3.6} = \frac{44}{36} \color{green}{\implies \frac{11}{9}}\]

OpenStudy (anonymous):

IS THAT ALL U DO ???

OpenStudy (anonymous):

I GET IT NOW THAT I LOOK AT IT AGAIN THANKS:)

OpenStudy (anonymous):

You want to do it more here? take a knife and kill your question then..